# On Mirzakhani’s twist torus conjecture

### v1

@@ 3.1. Twist tori
where the $\mathbb{R}^{n}$ fibers record the twisting about the curves of $\gamma$. In the case that $\gamma$ is a pair of pants, these are just the usual Fenchel–Nielsen coordinates. For an arbitrary measured lamination $\lambda$, the authors developed shear-shape coordinates that parametrize $\mathcal{T}_{g}$ in a similar way [CF21].
@@
In order to apply the results of [CF24] and connect hyperbolic twist tori to their flat counterparts, we will need to first lift twist tori to $\mathcal{P}^{1}\mathcal{M}_{g}$. To do so, choose any weight vector $\mathbf{h}=(h_{1},\ldots,h_{n})\in\mathbb{R}_{>0}^{n}$ satisfying $\bm{\ell}\cdot\mathbf{h}=1$. We can then lift the fiber $\text{cut}_{\gamma}^{-1}(Y)$ into $\mathcal{P}^{1}\mathcal{T}_{g}$ by setting the second coordinate equal to ${\mathbf{h}}\gamma$, the weighted multicurve ${h_{1}}\gamma_{1}\cup\ldots\cup{h_{n}}\gamma_{n}$. Quotienting out by the twist subgroup yields an embedded torus in $\mathcal{P}^{1}\mathcal{T}_{g}/\langle T_{\gamma_{i}}\rangle$ with its Lebesgue measure, and pushing down to moduli space we get a torus
@@ 3.2. Arc systems and ribbon graphs
Dually, the data of $\operatorname{\mathsf{Sp}}(Y)$ can also be recorded as a weighted, filling arc system on $Y$. Leaves of the orthogeodesic foliation $\mathcal{O}_{\partial Y}(Y)$ break up into finitely many isotopy classes of arcs rel $\partial Y$ and we use $\underline{\alpha}$ to denote the set of these isotopy classes. Assigning to each arc the length of its dual edge of $\operatorname{\mathsf{Sp}}(Y)$ then defines a positive weight system on $\underline{\alpha}$. See [CF21, §6] for a more thorough discussion of this duality.
@@
where $\mathcal{TRG}(S\setminus\gamma)$ is defined analogously to $\mathcal{T}(S\setminus\gamma)$; compare [AHC22, pg. 39]. We observe that when $\gamma$ is a pants decomposition, $\operatorname{\mathsf{Sp}}(\mathsf{Y})$ is completely determined by the boundary lengths $\bm{\ell}$. In the general case, the lengths of the $\gamma_{i}$ are explicitly obtainable from the metric on $\operatorname{\mathsf{Sp}}(\mathsf{Y})$, and we will continue to use $\bm{\ell}\in\mathbb{R}^{n}$ to denote the vector of lengths of the curves of $\gamma$.
@@ 3.3. Horospheres and slices
Following Forni [For21], we define a horospherical measure to be any probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h})$ that is absolutely continuous with respect to the Lebesgue class with continuous density and such that its restriction to each earthquake flow line is the restriction of Lebesgue to an interval. We also say that the pushforward of such a measure to $\mathcal{P}^{1}\mathcal{M}_{g}$ is horospherical.666 One could also consider any $\operatorname{Stab}(\mathbf{h}\gamma)$-invariant locally finite measure on $\mathcal{P}H_{\mathbf{h}\gamma}(L)$, take a local pushforward to the intermediate quotient $\mathcal{P}^{1}\mathcal{T}_{g}/\operatorname{Stab}(\mathbf{h}\gamma)$, and then take a further pushforward to $\mathcal{P}^{1}\mathcal{M}_{g}$, as in [AH21, Liu20, AHC22].
@@ 4.1. Twist tori and horospheres
The differentials $q\in\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ are completely horizontally periodic. With definitions as in [Wri15] or §5.1 below, the flat twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is just the orbit of any such by the abelian group $\mathbb{R}^{\gamma}$ of twists in the horizontal cylinders, and $\nu_{\gamma}(\mathsf{Y},\mathbf{h})$ is Lebesgue on $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$. We observe that $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is locally an affine submanifold in period coordinates, cut out by equations of the form $\operatorname{Im}(z)=h_{i}$ and $z=\ell(e)$, where $e$ is an edge of $\mathsf{Y}$ with length $\ell(e)$. By our convention that $\bm{\ell}\cdot\mathbf{h}=1$, the intersection of these affine subspaces lies in the unit area locus.
@@
which is actually a lift of the $1$-level set of the extremal length function $L_{\mathbf{h}\gamma}:\mathcal{T}_{g}\to\mathbb{R}_{>0}$ [CF21, GM91]. Using Theorem 2.1, $\mathcal{Q}H_{\gamma}(\mathbf{h})$ is a leaf of the strong unstable foliation. For any facet of $\mathcal{TRG}(S\setminus\gamma)$ we can also define the extremal length horosphere slice
@@
By construction of $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$ and Theorems 2.1 and 2.2, $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ is actually a leaf of the strong unstable foliation in the ambient stratum containing it. The angle around each singularity of the flat metric is $\pi$ times the valence of the corresponding vertex of $\Gamma$; see §2 of [CF21].
@@ 4.2. Affine invariant subvarieties
In general, the twist torus $\mathcal{Q}\mathbb{T}^{0}$ may not be minimal for the horocycle flow. In this case, we show that the $P$–orbit closure of a generic point (with respect to $\nu^{0}$) is $\mathcal{L}$; the rest of the argument then proceeds as above.
On any $q\in\mathcal{Q}\mathbb{T}^{0}$, choose a basis for relative homology consisting of horizontal saddle connections and saddles which connect the boundary components of the cylinders of $q$; this gives a local period coordinate chart for the stratum $\mathcal{Q}$. Fix a finite collection777 In fact, one can cover a lift of the twist torus to $\mathcal{Q}^{1}\mathcal{T}_{g}$ with a single chart of this form, but we use a finite collection of restricted charts so that we do not need to worry about the $\operatorname{Stab}(\gamma)$ action. of such charts that covers $\mathcal{Q}\mathbb{T}^{0}$. Let $\mathcal{N}\subsetneq\mathcal{L}$ be any affine invariant subvariety not containing $\mathcal{Q}\mathbb{T}^{0}$. In any of the period charts constructed above, $\mathcal{Q}\mathbb{T}^{0}$ is cut out by affine equations of the form $\operatorname{hol}(s)=\ell_{i}\in\mathbb{R}$ (specifying the length of the horizontal saddle connections) and equations of the form $\operatorname{Im}(\operatorname{hol}(s))=h_{i}$ (specifying the heights of the cylinders); compare §4.1. On the other hand, in period coordinates $\mathcal{N}$ is a union of finitely many linear subspaces, each of which is cut out by linear equations with real coefficients. Therefore, we see that in each chart, $\mathcal{Q}\mathbb{T}^{0}\cap\mathcal{N}$ is a union of finitely many proper subspaces, hence has $\nu^{0}$ measure 00. Moreover, we note that each subspace must be tangent to a proper subtorus of $\mathcal{Q}\mathbb{T}^{0}$, else its closure would be the entirety of $\mathcal{Q}\mathbb{T}^{0}$.
@@
is a union of countably many proper subtori of $\mathcal{Q}\mathbb{T}^{0}$. In particular, this set has measure zero with respect to $\nu^{0}$, and so the $P$–orbit closure of the generic point of $\mathcal{Q}\mathbb{T}^{0}$ is $\mathcal{L}$. ∎
@@ 4.3. Individual pushforwards
Forni also proved that if $\nu$ is especially nice, then the set $Z$ is empty. Recall that a horospherical measure $\nu$ is one that is supported in a leaf of the unstable foliation of $g_{t}$ on a stratum $\mathcal{Q}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$, is absolutely continuous with respect to the canonical affine measure on the leaf, has continuous density, and has conditional measures along horocycle orbits equal to the restriction of one-dimensional Lebesgue measures.888 Theorem 1.6 in Forni’s paper is stated for the stable foliation, but clearly applies to the unstable.
@@
From the definitions and the fact that $\mathcal{O}$ is a piecewise real-analytic homeomorphism on unstable leaves that takes earthquake flow to horocycle flow in a time-preserving fashion (Theorems 2.1 and 2.2), we see that if $\sigma$ is a horospherical probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h})$ (in the sense of §3.3), then $\mathcal{O}_{*}\sigma$ is horospherical with support $\mathcal{Q}H_{\gamma}(\mathbf{h})$ (in the sense above). Similarly, if $\sigma$ is a horospherical slice probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$, then $\mathcal{O}_{*}\sigma$ is supported on $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ and is horospherical for the ambient stratum component.
@@ 5.2. Lifting homologies
From this description, it follows that $\underline{\alpha}$ is co-orientable on $\Sigma$ if and only if the restriction of the horizontal foliation $\operatorname{Im}(q)$ to $\Sigma$ is orientable (see also the discussion in Section 7.2 of [CF21]), which happens if and only if the preimage $\widehat{\Sigma}$ is disconnected. Similarly, a joint orientation on $(\gamma,\underline{\alpha})$ is equivalent to a global choice of orientation of $\operatorname{Im}(q)$ on $S$, which implies the holonomy double cover is disconnected. ∎
@@ 5.3. Orbit closures of twist tori
Let $\gamma$ be any pants decomposition and fix any $\bm{\ell}$ and $\mathbf{h}$ as in the statement of the Theorem. Since pants are determined by their boundaries, $\bm{\ell}$ specifies a hyperbolic structure on $S\setminus\gamma$; let $\mathsf{Y}$ denote its spine and $\underline{\alpha}$ the dual arc system. We note that the edge lengths of $\mathsf{Y}$ are a linear function of the lengths $\bm{\ell}$ (see [CF21, Figure 1]), so it is equivalent to rescale $\bm{\ell}$ and to rescale $\mathsf{Y}$.

### v2

@@ 3.1. Twist tori
where the $\mathbb{R}^{n}$ fibers record the twisting about the curves of $\gamma$. In the case that $\gamma$ is a pair of pants, these are just the usual Fenchel–Nielsen coordinates. For an arbitrary measured lamination $\lambda$, the authors developed shear-shape coordinates that parametrize $\mathcal{T}_{g}$ in a similar way [CF24b].
@@
In order to apply the results of [CF24a] and connect hyperbolic twist tori to their flat counterparts, we will need to first lift twist tori to $\mathcal{P}^{1}\mathcal{M}_{g}$. To do so, choose any weight vector $\mathbf{h}=(h_{1},\ldots,h_{n})\in\mathbb{R}_{>0}^{n}$ satisfying $\bm{\ell}\cdot\mathbf{h}=1$. We can then lift the fiber $\text{cut}_{\gamma}^{-1}(Y)$ into $\mathcal{P}^{1}\mathcal{T}_{g}$ by setting the second coordinate equal to ${\mathbf{h}}\gamma$, the weighted multicurve ${h_{1}}\gamma_{1}\cup\ldots\cup{h_{n}}\gamma_{n}$. Quotienting out by the twist subgroup yields an embedded torus in $\mathcal{P}^{1}\mathcal{T}_{g}/\langle T_{\gamma_{i}}\rangle$ with its Lebesgue measure, and pushing down to moduli space we get a torus
@@ 3.2. Arc systems and ribbon graphs
Dually, the data of $\operatorname{\mathsf{Sp}}(Y)$ can also be recorded as a weighted, filling arc system on $Y$. Leaves of the orthogeodesic foliation $\mathcal{O}_{\partial Y}(Y)$ break up into finitely many isotopy classes of arcs rel $\partial Y$ and we use $\underline{\alpha}$ to denote the set of these isotopy classes. Assigning to each arc the length of its dual edge of $\operatorname{\mathsf{Sp}}(Y)$ then defines a positive weight system on $\underline{\alpha}$. See [CF24b, §6] for a more thorough discussion of this duality.
@@
where $\mathcal{TRG}(S\setminus\gamma)$ is defined analogously to $\mathcal{T}(S\setminus\gamma)$; compare [AHC22, pg. 39]. We observe that when $\gamma$ is a pants decomposition, $\operatorname{\mathsf{Sp}}(Y)$ is completely determined by the boundary lengths $\bm{\ell}$. In the general case, the lengths of the $\gamma_{i}$ are explicitly obtainable from the metric on $\operatorname{\mathsf{Sp}}(Y)$, and we will continue to use $\bm{\ell}\in\mathbb{R}^{n}$ to denote the vector of lengths of the curves of $\gamma$.
@@ 3.3. Horospheres and slices
Following Forni [For21], we define a horospherical measure to be any probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h})$ that is absolutely continuous with respect to the Lebesgue class with continuous density and such that its restriction to each earthquake flow line is the restriction of Lebesgue to an interval. We also say that the pushforward of such a measure to $\mathcal{P}^{1}\mathcal{M}_{g}$ is horospherical.111111 For example, in [AH21, Liu20, AHC22], such measures are obtained by taking certain $\operatorname{Stab}(\mathbf{h}\gamma)$-invariant measures on $\mathcal{P}H_{\gamma}(\mathbf{h})$, taking a local pushforward to the intermediate quotient $\mathcal{P}^{1}\mathcal{T}_{g}/\operatorname{Stab}(\mathbf{h}\gamma)$, then pushing down to $\mathcal{P}^{1}\mathcal{M}_{g}$.
@@ 4.1. Twist tori and horospheres
The differentials $q\in\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ are completely horizontally periodic, because their horizontal foliations are measure equivalent to $\mathbf{h}\gamma$. With definitions as in [Wri15] or §5.1 below, the flat twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is just the orbit of any such by the abelian group $\mathbb{R}^{\gamma}$ of twists in the horizontal cylinders, and $\nu_{\gamma}(\mathsf{Y},\mathbf{h})$ is Lebesgue on $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$. We observe that $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is locally an affine submanifold in period coordinates, cut out by equations of the form $\operatorname{Im}(z)=h_{i}$ and $z=\ell(e)$, where $e$ is an edge of $\mathsf{Y}$ with length $\ell(e)$. By our convention that $\bm{\ell}\cdot\mathbf{h}=1$, the intersection of these affine subspaces lies in the unit area locus.
@@
which is actually a lift of the $1$-level set of the extremal length function $L_{\mathbf{h}\gamma}:\mathcal{T}_{g}\to\mathbb{R}_{>0}$ [CF24b, GM91]. Using Theorem 2.1, $\mathcal{Q}H_{\gamma}(\mathbf{h})$ is a leaf of the strong unstable foliation. For any facet of $\mathcal{TRG}(S\setminus\gamma)$ we can also define the extremal length horosphere slice
@@
By construction of $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$ and Theorems 2.1 and 2.3, $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ is actually a leaf of the strong unstable foliation in the ambient stratum containing it. The angle around each singularity of the flat metric is $\pi$ times the valence of the corresponding vertex of $\Gamma$; see §2 of [CF24b].
@@ 4.2. Affine invariant subvarieties
In general, the twist torus $\mathcal{Q}\mathbb{T}^{0}$ may not be minimal for the horocycle flow. In this case, we show that the $P$–orbit closure of $\nu^{0}$–almost every point is $\mathcal{L}$; the rest of the argument then proceeds as above.
On any $q\in\mathcal{Q}\mathbb{T}^{0}$, choose a basis for relative homology consisting of horizontal saddle connections and saddles which connect the boundary components of the cylinders of $q$; this gives a local period coordinate chart for the stratum $\mathcal{Q}$. Fix a finite collection121212 In fact, one can cover a lift of the twist torus to $\mathcal{Q}^{1}\mathcal{T}_{g}$ with a single chart of this form, but we use a finite collection of restricted charts so that we do not need to worry about the $\operatorname{Stab}(\gamma)$ action. of such charts that covers $\mathcal{Q}\mathbb{T}^{0}$. Let $\mathcal{N}\subsetneq\mathcal{L}$ be any affine invariant subvariety not containing $\mathcal{Q}\mathbb{T}^{0}$. In any of the period charts constructed above, $\mathcal{Q}\mathbb{T}^{0}$ is cut out by affine equations of the form $\operatorname{hol}(s)=\ell_{i}\in\mathbb{R}$ (specifying the length of the horizontal saddle connections) and equations of the form $\operatorname{Im}(\operatorname{hol}(s))=h_{i}$ (specifying the heights of the cylinders); compare §4.1. On the other hand, in period coordinates $\mathcal{N}$ is a union of finitely many linear subspaces, each of which is cut out by linear equations with real coefficients. Therefore, we see that in each chart, $\mathcal{Q}\mathbb{T}^{0}\cap\mathcal{N}$ is a union of finitely many proper subspaces, hence has $\nu^{0}$ measure 00. Moreover, we note that each subspace must be tangent to a proper subtorus of $\mathcal{Q}\mathbb{T}^{0}$, else its closure would be the entirety of $\mathcal{Q}\mathbb{T}^{0}$.
@@
is a union of countably many proper subtori of $\mathcal{Q}\mathbb{T}^{0}$. In particular, this set has measure zero with respect to $\nu^{0}$, and so the $P$–orbit closure of almost every point of $\mathcal{Q}\mathbb{T}^{0}$ is $\mathcal{L}$. ∎
@@ 4.3. Individual pushforwards
Forni also proved that if $\nu$ is especially nice, then the set $Z$ is empty. Recall that a horospherical measure $\nu$ is one that is supported in a leaf of the unstable foliation of $g_{t}$ on a stratum $\mathcal{Q}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$, is absolutely continuous with respect to the canonical affine measure on the leaf, has continuous density, and has conditional measures along horocycle orbits equal to the restriction of one-dimensional Lebesgue measures.131313 Theorem 1.6 in Forni’s paper is stated for the stable foliation, but clearly applies to the unstable.
@@
From the definitions and the fact that $\mathcal{O}$ is a piecewise real-analytic homeomorphism on unstable leaves that takes earthquake flow to horocycle flow in a time-preserving fashion (Theorems 2.1 and 2.3), we see that if $\sigma$ is a horospherical probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h})$ (in the sense of §3.3), then $\mathcal{O}_{*}\sigma$ is horospherical with support $\mathcal{Q}H_{\gamma}(\mathbf{h})$ (in the sense above). Similarly, if $\sigma$ is a horospherical slice probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$, then $\mathcal{O}_{*}\sigma$ is supported on $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ and is horospherical for the ambient stratum component.
@@ 5.2. Lifting homologies
From this description, it follows that $\underline{\alpha}$ is co-orientable on $\Sigma$ if and only if the restriction of the horizontal foliation $\operatorname{Im}(q)$ to $\Sigma$ is orientable (see also the discussion in Section 7.2 of [CF24b]), which happens if and only if the preimage $\widehat{\Sigma}$ is disconnected. Similarly, a joint orientation on $(\gamma,\underline{\alpha})$ is equivalent to a global choice of orientation of $\operatorname{Im}(q)$ on $S$, which implies the holonomy double cover is disconnected. ∎
@@ 5.3. Orbit closures of twist tori
Let $\gamma$ be any pants decomposition and fix any $\bm{\ell}$ and $\mathbf{h}$ as in the statement of the Theorem. Since pants are determined by their boundaries, $\bm{\ell}$ specifies a hyperbolic structure on $S\setminus\gamma$; let $\mathsf{Y}$ denote its spine and $\underline{\alpha}$ the dual arc system. We note that the edge lengths of $\mathsf{Y}$ are a linear function of the lengths $\bm{\ell}$ (see [CF24b, Figure 1]), so it is equivalent to rescale $\bm{\ell}$ and to rescale $\mathsf{Y}$.